# Constrained Language Generation with Discrete Diffusion Models

### v1

@@ 4.2 Enforcing Constraints
where $\lambda$ is a Lagrange multiplier, and $\mu$ is a quadratic penalty term. When using a Lagrangian function, the optimization problem becomes
$\arg\min_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu),$
@@
and is a lower bound of the original projection operator (2). To obtain the strongest Lagrangian relaxation of the projection, the Lagrangian dual can be used to find the best Lagrangian multiplies $\lambda$ and penalty terms $\mu$, i.e.,
$\arg\max_{\lambda,\mu}\Bigl{(}\arg\min_{\bm{y}}\bigl{(}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)\bigr{)}\Bigr{)}.$ (3)
@@
For various classes of problems, the Lagrangian dual is a strong approximation of the original problem (our projection into the constraint space).
@@
2:Input: probability distribution $\mathcal{U}_{\theta}(\bm{x}_{t})$
3:$\bm{y}=\mathcal{U}_{\theta}(\bm{x}_{t})$
@@
5:     for $j=1$ to max_inner_iter do
6:         $\mathcal{L}_{KL}\leftarrow D_{\mathrm{KL}}(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y})$
7:         $\mathcal{L}_{\text{ALM}}\leftarrow\mathcal{L}_{KL}+\lambda\,\Delta g(\tilde{\phi}(\bm{y}))+\tfrac{\mu}{2}\,\Delta g(\tilde{\phi}(\bm{y}))^{2}$
8:         ${\bm{y}}\leftarrow{\bm{y}}-\eta\,\nabla_{\bm{y}}\mathcal{L}_{\text{ALM}}$
9:     end for
10:     $\lambda\leftarrow\lambda+\mu\,\Delta g(\bm{y}^{*})$
11:     $\mu\leftarrow\min\bigl{(}\alpha\mu,\,\mu_{\text{max}}\bigr{)}$
12:end while
13:$\bm{x}_{s}=\bm{y}$
14:return $\bm{x}_{s}$
@@
$\bm{y}$ $\leftarrow\bm{y}-\eta\nabla_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)$ (4a)
@@
where $\bm{y}^{\star}=\arg\max(\bm{y})$ is the discrete sequence implied by the distribution $\bm{y}$, $\eta$ is the gradient step size, $\alpha>1$ is a scaling factor that progressively increases $\mu$ over iterations, and $\mu_{\max}$ is an upper bound on the penalty term. These updates drive $\bm{y}$ as close as possible to satisfy $\Delta g(\tilde{\phi}(\bm{y}^{\star}))\leq\tau$ while also ensuring it remains close to the original denoised distribution ${\cal U}_{\theta}(\bm{x}_{t})$. We provide pseudocode for our implementation in Algorithm 1.
As we will show in the next section, practically these constraints can also be implemented as surrogate models (e.g., a classifier) that can be used to provide a continuous score, allowing for smooth gradient-base updates, as shown above. Additionally, we will also show that for token-level or logical constraints, such as enforcing the presence or absence of specific (combination of) tokens, $g$ can be an identity function, leading to sharper gradients and faster convergence in the iterative steps of Equation (4).
@@ 5 Experiments
To empirically demonstrate the advantage provided by CDD, we compare our method against similarly sized autoregressive and diffusion-based language models. Specifically, we benchmark our performance on constrained text generation against autoregressive baselines: GPT-2 and Llama 3.2 [41, 42], and discrete diffusion baselines MDLM and UDLM [16, 17]. We intentionally select models of similar size to, and up to one magnitude larger than, our discrete diffusion model, noting that we expect the performance of both classes of models to scale similarly with size. As shown by Schiff et al. [17], MDLM outperforms UDLM on the tasks described in Sections 5.1 and 5.2, and therefore we adopt this stronger baseline for comparison in these sections and utilize both in Section 5.3. For this reason, we use MDLM as the base discrete diffusion model for Sections 5.1 and 5.2, and UDLM for Section 5.3.
We evaluate on three constrained language generation tasks:
  1. 1.
    Toxicity Mitigation: Given a prompt from RealToxicityPrompts [43], the model is tasked with completing the passage with the goal of providing non-toxic responses.
  2. 2.
    Instruction Following: We present lexical constraints in an instruction-tuning format and ask the model to generate outputs that adhere to these specifications.
  3. 3.
    Molecule Generation: Our language model, trained on the molecular grammar ‘SMILES’ [44], is tasked to generate novel molecule sequences or sequences which satisfy certain properties.
5.1 Toxicity Mitigation
@@
While LLMs have demonstrated remarkable capabilities in generating human like text, a byproduct of data driven generation is the inadvertent creation of harmful, offense and dangerous text. Despite undergoing alignment finetuning, current state-of-the-art models still frequently generate harmful output, and while post-proccesing methods can be effective in moderating generations, they have still been shown to generate toxic output as these can be circumvented by adversarial or malicious prompts.
For this experiment, we generate output conditioned on prompts from the RealToxicityPrompts dataset [43]. We select prefixes with toxicity score above 0.5, in order to test how these models perform when given toxic inputs, and perplexity below 30.0, to avoid incoherent generations. To model the constraints, we finetune a GPT-Neo [45] model on the Jigsaw Toxic Comment Classification Challenge [^2] https://www.kaggle.com/c/jigsaw-toxic-comment- classification-challenge.
Figure 3 presents a comprehensive evaluation on model performance for toxicity mitigation, where the objective is to satisfy some toxicity threshold based on a surrogate model while maintaining competitive perplexity and language fluency. In this experiment, we assess several models architectures, baselines, and different toxicity thresholds. We report perplexity (PPL) at various percentiles (mean, 25th, 50th, and 75th), and violation rates at toxicity thresholds $\tau=0.25$, $\tau=0.50$, and $\tau=0.75$. All perplexity calculations are computed with respect to GPT2 (774M).
Among the baselines, GPT2 (124M) and Llama (1B) achieve the lowest perplexity score, however, they frequently produce toxicity content resulting in high violations. Although GPT2 + PPLM (345M) [30] is able to decrease the violations for all toxicity thresholds, it is unable to consistently prevent toxic generations and suffers from an increase in perplexity. MDLM (110M) exhibits higher perplexity than GPT2, with a median PPL of 39.75, and although toxicity violation rates are moderately improved relative to GPT2, they remain non-negligible. In contrast, our Constrained Discrete Diffusion (CDD) provides perfect constraint satisfaction at each toxicity thresholds, maintaining reasonable perplexity scores.
Toxicity Model Size PPL Viol (%)
Mean 25th 50th 75th $\tau=0.25$ $\tau=0.50$ $\tau=0.75$
GPT2 124M 19.143 14.57 17.75 21.61 36.3 23.8 16.2
@@
Llama 3.2 1B 15.66 11.52 14.58 17.95 34.9 27.8 23.1
MDLM 110M 46.72 26.57 39.75 57.55 32.1 23.2 17.2
$\textbf{CDD}_{\tau=0.25}$ (Ours) 110M 73.68 30.48 47.60 74.07 0.0 0.0 0.0
$\textbf{CDD}_{\tau=0.50}$ (Ours) 110M 60.49 29.14 45.64 68.10 0.0 0.0
$\textbf{CDD}_{\tau=0.75}$ (Ours) 110M 56.66 28.35 44.31 65.19 0.0
Figure 3: Results for various toxicity thresholds and PPL percentiles.
These results illustrate the trade-off between perplexity and toxicity mitigation. Baseline models such as GPT2 and Llama 3.2 achieve high fluency but at the cost of increased toxic outputs, whereas methods like GPT2 with PPLM provides neither strict avoidance of toxic generation nor the highest degrees of fluency. Our method, by contrast, provides a mechanism to establish a strict threshold on toxicity with respect to an external evaluator.
These findings are important as they show how, using projection-based control strategies, it is possible to achieve low toxicity levels at desired thresholds.
5.2 Instruction Following
Current state-of-the-art language models often struggle to satisfy seemingly simple generation requests from the user, and while instruction tuning has been shown effective in reducing the frequency that LLMs deviate from desired behaviors, these models still struggle to consistently satisfy user-defined constraints.
We evaluate CDD against such models using two types of constraints. First, we assess the ability of CDD to accurately answer how many letters are in a word, such as ‘How many R’s are in Strawberry?’. We additionally test lexical constraints based of the COLLIE dataset [46], prompting the models to generate word-indexed constrained sentences,
Model Size PPL Viol (%)
Counting GPT-2 (zero-shot) 124M 19.2 ± 0.8 100
GPT-2 (instruction-tuned) 124M 23.4 ± 1.3 5.0
Llama 3.2 (zero-shot) 1B 34.3 ± 3.8 100
Llama 3.2 (instruction-tuned) 1B 30.7 ± 2.9 9.1
MDLM 169M 16.9 ± 1.0 54.5
CDD (Ours) 169M 17.9 ± 1.1 0.0
Lexical GPT-2 (zero-shot) 124M 26.6 ± 4.2 97.6
GPT-2 (instruction-tuned) 124M 62.2 ± 2.9 96.4
Llama 3.2 (zero-shot) 1B 11.3 ± 2.0 97.2
Llama 3.2 (instruction-tuned) 1B 66.5 ± 6.1 93.6
MDLM 169M 21.9 ± 2.0 97.5
CDD (Ours) 169M 26.6 ± 4.1 0.0
Table 1: Results for functional constraints. PPL is assessed using GPT2 (774M). Bold and underlined values mark the best and second-best, respectively.
such as ‘Generate a sentence where the third-to-last word is mankind’. For each constraint we formulate a projection operator to impose hard constraints on the generation, pushing the samples back onto the constraint set.
5.2.1 Counting Constraints
As illustrated in the top of Table 1, unconstrained GPT-2 and Llama 3.2 fail to satisfy any letter-count requirement. Instruction tuning helps, reducing violation rates to 5.0% (GPT-2) and 9.1% (Llama 3.2), but still leaves a notable error margin. Although MDLM achieves the lowest perplexity, over half of its outputs violate the constraint. By contrast, CDD fully eliminates violations while preserving near-best fluency (17.9 perplexity), demonstrating that its constrained sampling process provides strict compliance without sacrificing language quality.
5.2.2 Lexical Constraints
Results are reported at the bottom of Table 1. Zero-shot GPT-2 and Llama 3.2 show similarly high violation rates (about 97%), and instruction tuning only minimally improves these figures. MDLM again delivers relatively low perplexity (21.9) but fails to address the precise positional requirements (97.5% violations). In stark contrast, CDD ensures perfect adherence, with only a moderate increase in perplexity, thus confirming the effectiveness of its projection mechanism for enforcing hard lexical constraints.
Taken together, these experiments reveal a fundamental limitation of conventional and instruction-tuned language models in scenarios requiring strict constraints. Conversely, CDD effectively produces outputs that satisfy both counting and lexical requirements by leveraging a projection operator to impose these constraints during generation.
Molecules (Synthetic Accessibility) Model Valid Novel QED Viol (%)
@@ 5.2.2 Lexical Constraints
AR 1023 11 0.46 91.6 78.4 62.3 42.0
$\text{AR}_{\text{FUDGE}\>\gamma=7}$ 925 13 0.48 11.1 9.8 9.6 9.2
MDLM 596 271 0.45 85.9 73.7 61.1 44.0
$\text{MDLM}_{\text{D-CFG}\>\gamma=3}$ 772 53 0.41 87.8 73.9 54.2 22.5
$\text{MDLM}_{\text{D-CBG}\>\gamma=3}$ 436 21 0.37 50.5 48.6 46.1 44.7
UDLM 895 345 0.47 89.4 88.0 58.1 37.8
$\text{UDLM}_{\text{D-CFG}\>\gamma=5}$ 850 69 0.47 80.6 58.6 35.9 13.9
$\text{UDLM}_{\text{D-CBG}\>\gamma=10}$ 896 374 0.47 90.1 77.8 58.6 37.7
$\textbf{CDD}_{\tau=3.0}$ (Ours) 353 36 0.63 0.0 0.0 0.0 0.0
$\textbf{CDD}_{\tau=3.5}$ (Ours) 863 91 0.62 - 0.0 0.0 0.0
$\textbf{CDD}_{\tau=4.0}$ (Ours) 936 108 0.61 - - 0.0 0.0
$\textbf{CDD}_{\tau=4.5}$ (Ours) 938 121 0.58 - - - 0.0
Figure 4: Results for synthetic accessibility constrained molecule generation constraints (left) and illustration of projection operation (right). QED and constraint violations are reported for only valid molecules. Bold and underlined values mark the best and second-best, respectively.
5.3 Molecular Generation
In this experiment we use UDLM to generate SMILES strings [44], a linear representation of a molecule’s structure with the vocabulary consisting of a finite set of symbols, and a grammar dictating molecular validity. Recent advances in generative modeling have enabled the design of molecular sequences by leveraging techniques from natural language processing. Despite their impressive ability to optimize for specific chemical properties, these models often generate molecules that fall short of practical requirements, either by producing compounds that are difficult to synthesize or by closely mimicking existing structures. Furthermore, generated molecules often fail to maximize qualitative properties such as drug-likeness (QED) [47] that are critical to the practicality of the generations and used as a qualitative metric for the experiments discussed in this section. This gap between in silico performance and laboratory feasibility has become a major hurdle in advancing molecular discovery.
To address this challenge, our framework incorporates two distinct constraint mechanisms. The first focuses on synthetic feasibility: by employing a surrogate scoring function, we ensure that every generated molecule meets a strict threshold for ease of synthesis, effectively prevent the generation of candidates that would be impractical to produce [48]. The second constraint targets molecular originality by rigorously projecting sequences that fall within a set of
Molecules (Novelty) Model Size Valid & Novel QED Viol (%)
@@ 5.3 Molecular Generation
$\text{AR}_{\text{D-CFG}\>\gamma=3}^{\dagger}$ 92M 79 0.60 91.61
$\text{MDLM}_{\text{D-CFG}\>\gamma=3}^{\dagger}$ 92M 96 0.60 69.82
$\text{UDLM}_{\text{D-CFG}\>\gamma=5}^{\dagger}$ 92M 64 0.62 93.69
$\textbf{CDD}_{\text{D-CFG}\>\gamma=5}(\text{Ours})$ 92M 251 0.60 0.00
@@
$\text{AR}_{\text{FUDGE}\>\gamma=7}^{\dagger}$ 92M 53 0.61 94.28
$\text{MDLM}_{\text{D-CBG}\>\gamma=3}^{\dagger}$ 92M 117 0.58 72.08
$\text{UDLM}_{\text{D-CBG}\>\gamma=10}^{\dagger}$ 92M 64 0.61 93.59
$\textbf{CDD}_{\text{D-CBG}\>\gamma=10}(\text{Ours})$ 92M 355 0.59 0.00
@@
Table 2: Results for novelty projection with and without QED guidance. Violation represents percentage of valid, but not novel molecule generations. QED is reported for only novel molecules. Results denoted with are as reported by Schiff et al. [17]. Bold and underlined values mark the best and second-best, respectively.
known compounds. For each constraint, we implement a dedicated projection operator that imposes hard limits on the generation process, resulting in final outputs that adhere precisely to our desired standards of synthetic accessibility or novelty. The ability to impose such constraints marks a crucial step toward the development of autonomous, LLM-driven laboratories for molecular innovation.
5.3.1 Synthetic Accessibility Constraint
While our molecular generation framework employs a surrogate model to optimize for synthetic accessibility during training, the actual assessment of accessibility violations is conducted by a separate, black-box external model. This external evaluation rigorously measures the degree to which generated molecules comply with the synthetic accessibility criteria, using a series of thresholds ($\tau=3.0$, $3.5$, $4.0$, and $4.5$). The results, summarized in Figure 4, confirm that CDD achieves perfect compliance across all thresholds. Across all thresholds, the CDD model yields a 0% violation rate, despite generating a competitive number of valid molecules and exhibiting the highest drug-likeness scores (referred to in the table as QED). These findings underscore the effectiveness of our hard constraint projection operators in ensuring that every candidate molecule not only scores well in property optimization but also adheres strictly to synthetic feasibility requirements as determined by an independent, external standard.
5.3.2 Novelty Constraint
Novelty in molecule generation refers to the model’s ability to produce new chemical structures that were not explicitly contained in its training set. In the context of generative modeling for drug design or other chemistry applications, novelty is often an important objective for the reasons of chemical space exploration and practical relevance. In real-world scenarios such as designing new drugs, researchers already have access to known molecules, thus, generating truly novel structures is crucial for discovering new compounds or structures.
In this experiment, we generate SMILES strings constraining all valid generations to be novel. We use baslines with no QED guidance alongside CFG and CBG to evaluate this methods impact on a desirable guidance improved metric. During the denoising step, if the candidate sequence is in an external database, we run a projection to generate the smallest possible change in probability space resulting in a novel molecule, effectively approximating the gradient step to this non-differentiable novelty indicator function. By applying this approach at every diffusion step, we ensure the model only produces new, previously unseen molecules, while simultaneously taking advantage of the creativity of the underlying discrete diffusion model.

### v2

@@ 4 Constraint Discrete Diffusion (CDD)
where $\lambda=(\lambda_{1},\ldots,\lambda_{m})$ is a non-negative Lagrange multiplier and $\mu=(\mu_{1},\ldots,\mu_{m})$ is a non-negative quadratic penalty term. When using a Lagrangian function, the primal optimization problem becomes
$\arg\min_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\bm{x}_{t}^{\prime};\lambda,\mu),$
@@
and is a lower bound of the original projection operator (5) by weak duality [41]. To obtain the strongest Lagrangian relaxation of the projection, the Lagrangian dual can be used to find the best Lagrangian multipliers $\lambda$ and penalty terms $\mu$, i.e.,
$\arg\max_{\lambda,\mu}\Bigl{(}\arg\min_{\bm{y}}\bigl{(}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\bm{x}_{t}^{\prime};\lambda,\mu)\bigr{)}\Bigr{)}.$ (7)
@@
In practice, the Lagrangian dual is a strong approximation of the original problem (our projection into the constraint space).
@@
2:Input: probability distribution $\bm{x}_{t}^{\prime}$
3:$\bm{y}=\bm{x}_{t}^{\prime}$
@@
5:     for $j\leftarrow 1$ to max_inner_iter do
6:         $\mathcal{L}_{KL}\leftarrow D_{\mathrm{KL}}(\bm{x}_{t}^{\prime}\|\bm{y})$
7:         $\mathcal{L}_{\text{viol}}\leftarrow\sum_{i=1}^{m}\lambda_{i}\,\Delta g_{i}(\tilde{\phi}(\bm{y}))+\tfrac{\mu_{i}}{2}\,\Delta g(\tilde{\phi}(\bm{y}))^{2}$
8:         $\mathcal{L}_{\text{ALM}}\leftarrow\mathcal{L}_{KL}+\mathcal{L}_{\text{viol}}$
9:         ${\bm{y}}\leftarrow{\bm{y}}-\eta\,\nabla_{\bm{y}}\mathcal{L}_{\text{ALM}}$
10:     end for
11:     $\lambda\leftarrow\lambda+\mu\,\Delta g(\bm{y}^{*})$
12:     $\mu\leftarrow\min\bigl{(}\alpha\mu,\,\mu_{\text{max}}\bigr{)}$
13:end while
14:$\bm{x}_{s}=\bm{y}$
15:return $\bm{x}_{s}$
@@
$\bm{y}$ $\leftarrow\bm{y}-\eta\nabla_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\bm{x}_{t}^{\prime};\lambda,\mu)$ (8a)
@@
where $\eta$ is the gradient step size, $\alpha>1$ is a scaling factor that progressively increases $\mu$ over iterations, and $\mu_{\max}$ is an upper bound on the penalty term. These updates drive $\bm{y}$ as close as possible to satisfy $\Delta g(\tilde{\phi}(\bm{y}^{\star}))\leq\tau$ while also ensuring it remains close to the original denoised distribution $\bm{x}_{t}^{\prime}$. Pseudocode is provided for our implementation in Algorithm 1.
To assess the robustness of this projection formulation, we performed an ablation study (Appendix D) examining the sensitivity of the Lagrangian hyper-parameters. Across the full grid of settings, the Lagrangian relaxation converges to feasible solutions in all our test cases.
As shown in the next section, there is a high degree of flexibility in how these constraints can be implemented. For instance, they can be implemented as surrogate models (e.g., a classifier) that can be used to provide a continuous score, allowing for smooth gradient-base updates, as shown above.
Theoretical justification.
The next result shows that the constrained reverse diffusion process converges to samples within the feasible region $\bm{C}$ while also keeping their distribution close to the data manifold, thus ensuring that the generated samples are both valid and realistic. Let $D_{\mathrm{KL}}(\bm{x}_{t},\bm{C})=\inf_{\bm{y}\in\bm{C}}D_{\mathrm{KL}}(\bm{y}\|\bm{x}_{t})$ denote the KL divergence from $\bm{x}_{t}$ to the set $\bm{C}$.
Theorem 4.1 (Convergence of CDD).
Let $\bm{C}$ be non-empty and $\beta$-prox-regular in the sense of [43, Def. 13.27], and the score network satisfy $\|\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})\|\leq G$ (a standard consequence of the bounded-data domain after normalization). Then, for positive step sizes $\gamma_{t},\leq\frac{1}{2G^{2}}\beta$, the following inequality holds for the distance to the feasible set $\bm{C}$:
$D_{\mathrm{KL}}(\bm{x}_{s}^{\prime},\bm{C})\;\leq\;(1-\bm{\upalpha}_{t})\,D_{\mathrm{KL}}(\bm{x}_{t}^{\prime},\bm{C})+\bm{\upalpha}_{t+1}^{2}G^{2},$
where $\bm{\upalpha}_{t}$ is proportional to the discrete Langevin step size $\gamma_{t}$ and $G$ bounds the score norm.
Theorem 4.1 shows that the distance to the feasible set $\bm{C}$ decreases at a rate of $1-\bm{\upalpha}_{t}$ at each step (up to an additive $\bm{\upalpha}_{t}^{2}G^{2}$ noise). This implies $\varepsilon$-feasibility after $\mathcal{O}(\bm{\upalpha}_{\min}^{-1})$ steps, with $\bm{\upalpha}_{\min}\propto\min_{t}\gamma_{t}$, and a cumulative KL drift within $\mathcal{O}(\sum_{t}\bm{\upalpha}_{t})$. Missing proofs are provided in Appendix F.
@@ 5 Experiments
To empirically demonstrate the advantage provided by CDD, we compare our method against similarly sized autoregressive and diffusion-based language models. Specifically, we benchmark our performance on constrained text generation against autoregressive baselines: GPT-2 and Llama 3.2 [44, 45], and discrete diffusion baselines MDLM and UDLM [7, 8]. We intentionally select models of similar size to, and up to one magnitude larger than, our base discrete diffusion model, noting that we expect the performance of both classes of models to scale similarly with size. As demonstrated by Schiff et al. [8], MDLM outperforms UDLM on the natural language tasks discussed in Sections 5.1 and 5.3. Consequently, we compare to both models in Section 5.2 and solely MDLM for the following natural language tasks. We use UDLM as the base discrete diffusion model for Section 5.2, and MDLM in Sections 5.1 and 5.3. For each application, CDD uses configurations as described in [7, 8] unless otherwise specified. Additional experiment details are available in Appendix  C.
5.1 Natural Language Toxicity Mitigation
@@
While LLMs have demonstrated remarkable capabilities in generating human like text, a byproduct of data driven generation is the inadvertent creation of harmful, offensive and dangerous text. Despite undergoing alignment finetuning, current state-of-the-art models still frequently generate harmful outputs. While post-processing methods can be effective in moderating generations, they can be circumvented by adversarial or malicious prompts, resulting in toxic output generation. In this experiment, we generate outputs conditioned on prompts from the RealToxicityPrompts dataset [46]. We select prefixes with toxicity score above 0.5, testing how these models perform on toxic inputs, and perplexity below 30.0, avoiding incoherent generations. To model the constraints, GPT-Neo [47] is finetuned on the Jigsaw Toxic Comment Classification Challenge [^2] https://www.kaggle.com/c/jigsaw-toxic-comment- classification-challenge. We evaluate on 1,000 samples, and discuss ALM hyper-parameters in Appendix  D, and runtime in Appendix  E.
Toxicity
Model Size PPL LLM-Judge Coherence Viol (%)
Mean Median Coherence Change (%) $\tau=0.25$ $\tau=0.50$ $\tau=0.75$
GPT2 124M 18.78 17.96 42.68 33.2 21.6 13.1
@@
[…]
Llama 3.2 1B 15.66 14.58 57.10 34.9 27.8 23.1
MDLM 110M 46.72 39.75 20.02 32.1 23.2 17.2
$\textbf{CDD}_{\tau=0.25}$ (Ours) 110M 61.55 45.42 20.16 +0.7 0.0 0.0 0.0
$\textbf{CDD}_{\tau=0.50}$ (Ours) 110M 59.44 44.24 20.30 +1.4 0.0 0.0
$\textbf{CDD}_{\tau=0.75}$ (Ours) 110M 54.87 43.51 20.88 +4.3 0.0
Figure 3: Results across different toxicity thresholds, PPL percentiles, LLM-Judge coherence, and coherence degradation levels. PPL is evaluated with GPT-2-XL, and coherence metrics are assessed by the LLM-Judge. Bold and underlined values denote the best and second-best, respectively.
Figure 3 presents a comprehensive evaluation on model performance for toxicity mitigation, where the objective is to have the generated output satisfy some toxicity threshold, based on a surrogate model, while maintaining competitive perplexity and coherence. We report perplexity (PPL), coherence as determined by a LLM judge, and violation rates at toxicity thresholds $\tau=0.25$, $\tau=0.50$, and $\tau=0.75$ across different model architecures and baselines.
Among the baselines, GPT2 (124M) and Llama (1B) achieve the lowest perplexity and highest coherence score, however, they frequently produce toxicity content resulting in high violations. Although PPLM and FUDGE (124M) [26, 15] are able to decrease the violations for all toxicity thresholds, they are unable to consistently prevent toxic generations while suffering from an increase in perplexity and a decrease in coherence. MDLM (110M) exhibits higher perplexity than GPT2 and lower coherence, and similarly shares a non-negligible amount of toxicity constraint violations. In contrast, CDD provides perfect constraint satisfaction at each toxicity thresholds, maintaining comparable perplexity scores to its base diffusion language model and other baselines. Additionally, while constraint imposition degrades the coherence of the controllable generation baselines, CDD observes an increase in coherence as evaluated by the the LLM-Judge. These findings are important as they show how, using projection-based control strategies, it is possible to achieve complete toxicity constraint satisfaction at desired thresholds.
5.2 Molecular Generation
Next, we show the ability of CDD to impose constraints on a scientific task, specifically in the domain of molecular generation. In this experiment we generate SMILES strings [48], a linear representation of a molecule’s structure with a vocabulary consisting of a limited vocabulary of forty symbols and a grammar dictating molecular validity. Recent advances in generative modeling have enabled the design of molecular sequences by leveraging techniques from natural language processing. Despite their impressive ability to optimize for specific chemical properties, these models often generate molecules that fall short of practical requirements, either by producing compounds that are difficult to synthesize or by closely mimicking existing structures. Furthermore, generated molecules often fail to maximize qualitative properties such as drug-likeness (QED) [49] that are critical to the practicality of the generations and used as a qualitative metric for the experiments discussed in this section.
To address this challenge, our framework incorporates two distinct constraint mechanisms: (1) synthetic feasibility, which ensures that generated molecules can be synthesized in a laboratory setting [50], and (2) novelty, which guarantees that the generated molecules are not already present in training datasets. For each constraint, we implement a dedicated projection operator, described in Appendix C, that imposes hard limits on the generation process, resulting in final outputs that adhere precisely to our desired standards of synthetic accessibility or novelty.
Synthetic accessibility constraints. Synthetic accessibility is commonly assessed using black-box, non-differentiable functions. To enable gradient-based optimization, we train a surrogate model on the QM9 dataset [51], where training labels are derived from RDKit’s synthetic accessibility score [52]. This surrogate acts as our measurement function $g(\cdot)$ that approximates the original non-differentiable synthetic accessibility scorer, allowing construction of the projection operator depicted in Figure 4. While our molecular generation framework employs this surrogate model to optimize for synthetic accessibility during training, the actual assessment of accessibility violations is conducted by a separate, black-box external model. This external evaluation rigorously measures the degree to which generated molecules comply with the synthetic accessibility criteria, using a series of thresholds ($\tau=3.0$, $3.5$, $4.0$, and $4.5$). For the discrete diffusion baselines, we adapt the guidance mechanisms provided by [8] for QED by retraining on synthetic accessibility labels for Classifier-Based Guidance (CBG) and Classifier-Free Guidance (CFG). A similar adaptation is applied for the autoregressive baselines employing FUDGE guidance [15]. The results, summarized in Figure 5 (left), confirm that CDD achieves perfect compliance across all thresholds. This 0% violation rate is achieved, all while generating a competitive number of valid molecules and exhibiting the highest drug-likeness scores.
Figure 4: SA projection illustration.
Molecules (Synthetic Accessibility)
Model Valid Novel QED Viol (%)
@@ 5.2 Molecular Generation
AR 1023 0 0.46 91.6 78.4 62.3 42.0
$\text{AR}_{\text{FUDGE}\>\gamma=7}$ 925 6 0.48 11.1 9.8 9.6 9.2
MDLM 596 20 0.45 85.9 73.7 61.1 44.0
$\text{MDLM}_{\text{D-CFG}\>\gamma=3}$ 772 3 0.41 87.8 73.9 54.2 22.5
$\text{MDLM}_{\text{D-CBG}\>\gamma=3}$ 436 14 0.37 50.5 48.6 46.1 44.7
UDLM 895 21 0.47 89.4 88.0 58.1 37.8
$\text{UDLM}_{\text{D-CFG}\>\gamma=5}$ 850 18 0.47 80.6 58.6 35.9 13.9
$\text{UDLM}_{\text{D-CBG}\>\gamma=10}$ 896 28 0.47 90.1 77.8 58.6 37.7
[…]
Molecules (Novelty)
Model Size Valid & Novel QED Viol (%)
@@
ARD-CFG $\gamma=3$$\dagger$ 92M 79 0.60 91.61
MDLMD-CFG $\gamma=3$$\dagger$ 92M 96 0.60 69.82
UDLMD-CFG $\gamma=5$$\dagger$ 92M 64 0.62 93.69
CDDD-CFG $\gamma=5$ (Ours) 92M 251 0.60 0.00
@@
ARFUDGE $\gamma=7$$\dagger$ 92M 53 0.61 94.28
MDLMD-CBG $\gamma=3$$\dagger$ 92M 117 0.58 72.08
UDLMD-CBG $\gamma=10$$\dagger$ 92M 64 0.61 93.59
CDDD-CBG $\gamma=10$ (Ours) 92M 355 0.59 0.00
@@
Figure 5: Left: Results for synthetic accessibility constrained molecule generation constraints. QED and constraint violations are reported for only valid molecules, and novel molecules must be valid and have no violation ($\tau\leq 3.0$). Right: Results for novelty projection with and without QED guidance. Violation represents percentage of valid, but not novel molecule generations. QED is reported for only novel molecules. Results denoted with are as reported by Schiff et al. [8]. Bold and underlined values mark the best and second-best, respectively.
Novelty constraints. Novelty in molecule generation refers to the model’s ability to produce new chemical structures that were not explicitly contained in its training set. In the context of generative modeling for drug design or other chemistry applications, novelty is often an important objective for the reasons of chemical space exploration and practical relevance. In this experiment, we generate SMILES strings constraining all valid generations to be novel. The novelty constraint is enforced at every denoising step. If the current candidate sequence already exists in an external database, a projection operator minimally perturbs its token-probability vector to yield an unseen molecule, thereby approximating a gradient step through the otherwise non-differentiable novelty indicator. Specifically, the novelty projection operator is a best-first traversal of the token-probability space: each token flip incurs a cost equal to its probability gap from the argmax, a priority queue retrieves the lowest-cost unseen sequence, and we then renormalize and cache it to prevent duplication. As it selects the sequence with the minimal cumulative flip cost, our distance metric over sequences, the procedure is mathematically equivalent to projecting the distribution onto the set of novel sequences. Repeating this operation throughout the diffusion trajectory guarantees that only new compounds are emitted while preserving the exploratory capacity of the discrete diffusion model.